1. In which case is the bond polarity incorrect?
a. δ+H–Fδ–
b. δ+Si–Sδ–
c. δ+Cl–Iδ–
d. δ+K–Oδ–
e. δ+Mg–Hδ–
2. Calculate the value of ΔH° for the reaction C2H6 + 2 Cl2 → C2H4Cl2 + 2 HCl given the bond energies (kJ/mol):

H-H 432 kJ/mol

C-H 314 kJ/mol

H-Cl 427 kJ/mol

C-C 347 kJ/mol

Cl-Cl 243 kJ/mol

C-Cl 339 kJ/mol
3. What intermolecular forces are responsible for base pairing in DNA?

a. Hydrogen bonding
b. London
c. Hydrophobic
d. Dipole-dipole
1
Expert's answer
2019-11-22T04:34:10-0500
  1. C
  2. ΔH°= (314*2 + 243*2) - (339*2 + 427*2) = -418 (kJ/mol)
  3. A